Quantity of shoes is s and calculators be c, total income of the consumers be M. So, ms/2 + mc/2 = m. This implies s+c =2 Let the total number of labour be L.

labour in Italy = 80 Labour in Sweden = 60

Slope of PPC for Italy = -2/1 = -2 = -P​​​​​c​​​​/P​​​​​​s = P

Slope of PPC for Sweden = -2/4 = -1/2= -P​​​​​c​​​​/P​​​​​​w = P*

so Italy has a comparative advantage in producing Calculators and Sweden has a comparative advantage in producing shoes.

Italy

Sweden

Shoes (s)

1

4

Calculators (c)

2

2

Question

d) Under free trade, which country produces which good(s)? How many units?

e) Who gains from trade? Who loses from trade? State labours’ stance towards free trade in each country.
Answer #1

d) Under free trade, Italy will produce calculators and Sweden will produce shoes. Italy will produce c units of calculators and Sweden will produce s units of shoes. As Italy has 80 labors, number of calculators produced = 80*2 = 160. As Sweden has 60 labors, number of shoes produced=60*4 = 240.

e) To gain from trade, world price must be in the range of $1/2 =$0.5 to $2 (Opportunity costs for both countries) . Let the world price be $1 , i.e, 1 shoe will be exchanged for 1 calculator. As Sweden produces more units of shoes (240) than Italy, which produces less number of calculators, it gains from trade and Italy loses from trade. This might result in relocation of some Italian workers to Sweden.